OpenStudy (anonymous):

4(2x -3) + 8(x - 4) = 2(2x + 6)

OpenStudy (anonymous):

first distribute the 4, 8, and 2

undeadknight26 (undeadknight26):

What they said :D ^

OpenStudy (anonymous):

8x-12+8x-32=4x+12??

OpenStudy (anonymous):

yep

undeadknight26 (undeadknight26):

I believe thats right...bit rusty on this stuff XD

OpenStudy (anonymous):

now combine like terms

OpenStudy (anonymous):

16x-44=4x+12?

OpenStudy (anonymous):

yep :) now try to get both x's on the same side. you can either subtract 16 from both sides, or subtract 4 from both sides

OpenStudy (anonymous):

12x-44=12?

OpenStudy (anonymous):

now get x by itself

OpenStudy (anonymous):

which I add 44 to the 12 by itself correct?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

so then you have 12x=56

OpenStudy (anonymous):

answer is 56/12 right?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

Thanks so much! :)

OpenStudy (anonymous):

mhm :)

OpenStudy (anonymous):

reduces to 14/3 btw

OpenStudy (anonymous):

thanks :)